Use the following information about Company X to help answer questions 14-20:

\ start of 2017. :  Issued 200M shares of stock at $25 per share $500M  Issued $400M of long term debt; as part of this interest only loan, the company agreed to pay creditors 15.0% per year; this implies the current portion of long term debt is $0 in 2017/2018

At the end of 2017  $400M in sales, $260M in operating earnings, and $160 in net income  $500M Gross PPE , $450 Net PPE  Working Capital of $430, excluding marketable securities  The company paid a dividend of 30 cents per share  Investors were pleased with the results; Company X’s share price closed at $28 per share

At the end of 2018:  $480M in sales, $300M in operating earnings, and $200 in net income  $600M Gross PPE , $490 Net PPE  Working Capital of $460, excluding marketable securities  The company paid a dividend of 40 cents per share  Investors were pleased with the results; Company X’s share price closed at $33 per share

Throughout 2017 and 2018:  The company did not issue or repurchase any shares of stock other than during the IPO  The company did not issue or retire any long term debt other than during the initial debt offering

18. Do not include marketable securities or the current portion of long term debt when calculating working capital investment. In 2018, Company X’s Cash Flow from Operations was closest to:

a. $180

b. $220

c. $230

d. $280

e. $290

19. Ignore your answer to question 18 and assume that CFO was $250. Calculate fixed capital investment as the year on year change in Gross PPE. In 2018, Company X’s Free cash Flow to Firm was closest to:

a. $110

b. $150

c. $170

d. $190

e. $210

20. In 2017, Company X’s Return on Equity, where ROE=NI/EOY SE, is closest to:

a. 20%

b. 23%

c. 25%

d. 27%

e. 33%
Answer #1

Question 18:

Free cash flow for 2018 = operating Income + Depreciation - Increase in working capital - Capital expenditure

Operating income (2018) = 300 M

Accumulated depreciation in 2017 = 500-450 = 50

Accumulated depreciation in 2018 = 600-490 = 110

Depreciation in 2018 = 110-50 =60

Deprecation (2018) = 60 M

Increase in working capital = Working capital in 2018 - Working capital in 2017

Increase in working capital = 460-430

Increase in working capital (2018) = 30 M

Capital expenditure = Net PPE 2018 - Net PPE 2017

Capital expenditure = 490-450

Capital expenditure (2018) = 40 M

Free cash flow = 300 + 60 -30 -40 = 290 M (Option D)

Note: We have answered the first full Multiple choice question. Please not that only one multiple choice question can be answered at a time. Please post other separately for experts to answer